OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 7 tan3x - 21 tan x = 0

hartnn (hartnn):

its \(\tan 3x\) or \(\tan^3x\) ??

OpenStudy (anonymous):

tan^3 sorry

hartnn (hartnn):

okk, factor out 7tan x from left side what u get ?

OpenStudy (anonymous):

That mainly my confusion, I missed a day and I don't know how to factor out 7tan x. How do I factor?

hartnn (hartnn):

factoring 'x' from \(x^3\) gives \(x (x^2)\) right ?

hartnn (hartnn):

so, factoring tan x from \(\tan^3x\) will give you \(\tan x (\tan^2x)\) so your left side becomes \(\Large 7 \tan x (\tan^2 x-3) = 0 \) see if you get this step properly...

OpenStudy (anonymous):

oh okay I got that. How do I find the solutions after factoring

hartnn (hartnn):

when ab = 0 a= 0 or b = 0 so we get \( \tan x = 0\) or \(\tan^2x = 3\) can you find for which angles are the above equations true ?

OpenStudy (anonymous):

\[\tan ^{2} x=3\] ?

hartnn (hartnn):

tan x = 0 for which angles in 0 to 2pi ?? tan^2 x = 3 so, \( \tan x= \sqrt3\) for which angles in 0 to 2pi ?? or \( \tan x= -\sqrt3\) for which angles in 0 to 2pi ??

OpenStudy (anonymous):

tan x= -\[\sqrt{3}\]

hartnn (hartnn):

for what values of x??

OpenStudy (anonymous):

am i finding out x?

hartnn (hartnn):

yes...

OpenStudy (anonymous):

Okay. idk how to find x

hartnn (hartnn):

do you have unit circle ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

good, for tan x to be -sqrt 3 sin x = -sqrt 3/2 cos x = 1/2 for which angle ? or sin x = sqrt 3/2 cos x = -1/2 for which angle ?

OpenStudy (anonymous):

I'm not understanding this

hartnn (hartnn):

didn't you attempt easier problems before going to this one? like sin x =0 or cos x = 1/2 something like that ? if not, this is a difficult problem to start with...

OpenStudy (anonymous):

No he just gave me this one and told me to solve it...

OpenStudy (anonymous):

Can you show me the steps in finding the answer?

hartnn (hartnn):

can you use calculator ?

OpenStudy (anonymous):

Yes I can

hartnn (hartnn):

cool, so can't you just plug in tan x = -sqrt 3 in you calculator and it gives you x values ??

OpenStudy (anonymous):

Its shoes me its graph of tan x =-3^2